# Beyond Cosine Decay: On the effectiveness of Infinite Learning Rate Schedule for Continual Pre-training

### v2

@@ 5.2 Results for pre-training mae on multiple datasets
We begin pre-training on the DCLM dataset and observe that even in this pre-training phase, rapid annealing in the case of infinite schedule yields a lower validation loss compared to the cosine schedule, offering a competitive advantage. This trend is evident in Figure 7, with further details provided in Appendix D. raining on DCLM, we continue training on the Stack dataset. Figure 3 shows the validation loss on the DCLM ($\mathcal{D}_{0}$) and Stack ($\mathcal{D}_{1}$) dataset for cosine and infinite schedule with varying $\eta_{const}$ and $P$. We observe that all the configurations of infinite schedule helps in mitigating catastrophic forgetting with a lower validation loss on DCLM data, as compared to cosine, with a minimum validation loss for $\eta_{const}=1e-4$ and longer cooldown of $P=0.6$. This is in concurrence with the observations for MAE large scale pre-training.
@@
To alleviate forgetting, we further introduce a replay mechanism where we sample $50\%$ of the data from the previous task (DCLM) and $50\%$ from the current task (Stack). Figure 4 shows the validation loss on the DCLM ($\mathcal{D}_{0}$) and Stack ($\mathcal{D}_{1}$) dataset for cosine and infinite schedule with varying $\eta_{const}$ and $P$ with replay. We observe that the infinite schedule with $\eta_{const}=2e-4$ and longer cooldown of $P=0.6$ helps in mitigating catastrophic forgetting with minimum validation loss, as compared to cosine and other configurations of infinite scheduling. We further observe that infinite schedule, irrespective of the $P$ and $\eta_{const}$ gives a lower validation loss as compared to cosine. A higher $\eta_{const}$ likely enhances adaptability to the current task, while a lower $\eta_{const}$ minimizes forgetting on previous tasks. Since replay mitigates forgetting, a higher $\eta_{const}$ ultimately achieves the best overall performance, balancing adaptability and retention.
To further strengthen our evaluation, we introduce a language shift by continually pre-training on the German dataset (German language). This transition imposes a more pronounced distributional shift, as the model moves from programming language data (Stack) to natural language. As in previous sections, we measure validation loss across all datasets while continually pre-training on German without replay. Given our earlier findings that short cooldown proportions are detrimental, we train models only with $P=0.6$ under an infinite schedule. Consistent with our previous observations (Figure 3), we find that the infinite schedule with $\eta_{const}=1e-4$ and $P=0.6$ yields the best performance in mitigating forgetting.
@@
Cosine 100B DCLM $\rightarrow$ 100B Stack 33.17 31.79 25.2 49.17 42.72 62.51 25.49 38.58
@@
100B DCLM $\rightarrow$ 100B Stack (50% Replay) 49.16 43.72 32.6 52.09 50.59 68.93 27.65 46.39
@@
Table 3: Zero-shot results on popular LM benchmarks. Normalized accuracy is reported. We observe on average, as expected, that the infinite schedule with $\eta_{const}=2e-4$ with a $50\%$ replay gives the best performance across all the benchmarks. Even without Replay, both the infinite schedules give better performance as compared to cosine. This demonstrates the effectiveness of infinite schedule in mitigating forgetting.
@@ 7 Conclusion
Our results demonstrate that infinite cosine schedules effectively reduce catastrophic forgetting in CPT of foundation models across diverse domains, thus proving to be a robust and scalable improvement over repeated cosine annealing. We saw that infinite schedules allow us to seamlessly resume training (continued pre-training at $\eta_{const}$), that alone or along with replay, they best continual learning baselines, and, on large-scale experiments across multiple vision and language datasets, they consistently outperform repeated cosine decay.
@@
We acknowledge support from NSERC Discovery Grant RGPIN- 2021-04104 [E.B.], FRQNT New Scholar [E.B.], the Canada CIFAR AI Chair Program [I.R.], and the Canada Excellence Research Chairs Program [I.R.]. We would also like to acknowledge funding from the FRQNT Doctoral (B2X) scholarship [B.T.]. This research was made possible thanks to the computing resources on the Frontier supercomputer, provided as a part of the ALCC 2024 program award “Scalable Foundation Models for Transferable Generalist AI”. These resources were provided by the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. In particular, we thank Jens Glaser for his help with the Summit supercomputer.
@@ A.2 Hyperparameters and Implementation details for CIFAR10 MAE
For our architecture, we employ a ViT-tiny encoder (12 layers, 192 hidden dimension, 3 attention heads) to match the scale of CIFAR-10, with our implementation based on the Zhang (2021)’s work. Our model uses a masking ratio of 0.75, consistent with the original MAE, and incorporates a lightweight decoder (4 layers) with learned position embeddings to reconstruct the masked patches. Regarding the learning rate configuration, we selected a maximum learning rate of 7.5e-5 through hyperparameter tuning over the values [7e-5, 1.5e-4, 3e-4] on the first two tasks, with a minimum learning rate of 7.5e-6. For most experiments, we employ a constant learning rate of 1.875e-5 and a cooldown proportion of 0.4, except for experiments without replay where we increase the constant learning rate to 5.625e-5. These optimal values were determined through experiments similar to our large-scale setup, testing cooldown proportions [0.3, 0.4, 0.5] and constant learning rates [1.875e-5, 5.625e-5]. While these findings align with our large-scale experiments, the small dataset size necessitated a slightly larger cooldown proportion to maintain a higher learning rate for a longer duration. For linear probing experiments in our small-scale setup, we utilized the AdamW(Loshchilov & Hutter, 2019) optimizer with a weight decay coefficient of $35e-$ 3 and momentum parameters $\beta_{1}$ and $\beta_{2}$ set to 0.9 and 0.95 respectively. The linear probing experiments implemented a cosine decay learning rate schedule with a maximum learning rate $\eta_{max}=1e-3$, running for 100 epochs total, including 10 warmup epochs, with a batch size of 128. Complete hyperparameter details for pre-training and linear probing can be found in the corresponding tables Table 6 and Table 6. For the baseline methods MAS (Aljundi et al., 2018) and LwF (Li & Hoiem, 2017), we conducted hyperparameter tuning using grid search over the first two tasks. For MAS, we explored values of $\alpha$ and $\lambda$ in [0.25, 0.5, 0.75]. Similarly for LwF, we searched for optimal $\alpha$ values within the same range.
@@ Appendix C Implementation details and hyperparameters for language pre-training
Warmup percent ($N_{\textit{w}}$) $1$
@@
Warmup percent ($N_{\textit{w}}$) $1$
Cooldown iters percent ($N_{\textit{c}}$) 00
@@ Appendix D Pretraining with DCLM data
Figure 7 shows the validation loss on the DCLM dataset for cosine and infinite schedule with varying $\eta_{const}$ and cooldown proportion $P$. We observe that the infinite schedule with a higher constant learning rate ($\eta_{const}=2e-4$) and cooldown proportion ($P=0.6$) performs better than the cosine schedule and the other configurations of the infinite schedule. The final checkpoint, in the case of infinite schedule, is obtained via annealing which we perform for $15\%$ of the total iterations after the constant phase, as shown in Figure 1. It can be inferred that the infinite schedule with $\eta_{const}=2e-5$ and $P=0.6$ performs the best, with validation loss rapidly decaying in the annealing phase. We also observe that a shorter cooldown phase ($P=0.3$) results in suboptimal performance with higher validation loss, thus indicating that a longer cooldown phase is beneficial. We note that this corresponds to $28K$ steps. As for the $\eta_{const}$, we observe that both $41e-$ 4 and $42e-$ 4 perform similarly, with the latter having a slightly lower validation loss, indicating that a higher constant learning rate gives a better exploration possibility during training.

### v3

@@ 5.3 Results for Continual Pre-training LLMs
We begin pre-training on the DCLM dataset and observe that even in this pre-training phase, rapid annealing in the case of infinite schedule yields a lower validation loss compared to the cosine schedule, offering a competitive advantage. This trend is evident in Figure 7, with further details provided in Appendix D. After pre-training on DCLM, we continue training on the Stack dataset. Figure 3 shows the validation loss on the DCLM ($\mathcal{D}_{0}$) and Stack ($\mathcal{D}_{1}$) dataset for cosine and infinite schedule with varying $\eta_{const}$ and $P$. We observe that all the configurations of infinite schedule helps in mitigating catastrophic forgetting with a lower validation loss on DCLM data, as compared to cosine, with a minimum validation loss for $\eta_{const}=1e-4$ and longer cooldown of $P=0.6$. This is in concurrence with the observations for MAE large scale pre-training.
@@
To alleviate forgetting, we further introduce a replay mechanism where we sample $50\%$ of the data from the previous task (DCLM) and $50\%$ from the current task (Stack). Figure Figure 4 shows the validation loss on the DCLM ($\mathcal{D}_{0}$) and Stack ($\mathcal{D}_{1}$) dataset for cosine and infinite schedule with varying $\eta_{const}$ and $P$ with replay. We observe that the infinite schedule with $\eta_{const}=2e-4$ and longer cooldown of $P=0.6$ helps in mitigating catastrophic forgetting with minimum validation loss, as compared to cosine and other configurations of infinite scheduling. We further observe that infinite schedule, irrespective of the $P$ and $\eta_{const}$ gives a lower validation loss as compared to cosine. A higher $\eta_{const}$ likely enhances adaptability to the current task, while a lower $\eta_{const}$ minimizes forgetting on previous tasks. Since replay mitigates forgetting, a higher $\eta_{const}$ ultimately achieves the best overall performance, balancing adaptability and retention. Furthermore, we present experiments in Appendix F that demonstrate the flexibility and agility of the infinite cosine schedule in preserving knowledge from previous tasks.
To further strengthen our evaluation, we introduce a language shift by continually pre-training on the German dataset (German language). This transition imposes a more pronounced distributional shift, as the model moves from programming language data (Stack) to natural language. As in previous sections, we measure validation loss across all datasets while continually pre-training on German as shown in Figure 5 Given our earlier findings that short cooldown proportions are detrimental, we train models only with $P=0.6$ under an infinite schedule. Consistent with our previous observations (Figure 3), we find that the infinite schedule with $\eta_{const}=1e-4$ and $P=0.6$ yields the best performance in mitigating forgetting.
@@
Cosine 100B DCLM $\cup$ 100B Stack 46.36 42.71 32.0 50.28 51.05 68.44 27.50 45.47
100B DCLM $\rightarrow$ 100B Stack 33.17 31.79 25.2 49.17 42.72 62.51 25.49 38.58
@@
100B DCLM $\rightarrow$ 100B Stack $\rightarrow$ 100B German 18.32 29.74 23.20 50.90 33.83 58.92 28.41 33.92
100B DCLM $\rightarrow$ 100B Stack $\rightarrow$ 100B German (50% Replay) 44.09 40.31 30.4 53.74 48.69 67.51 27.80 44.65
@@
100B DCLM $\rightarrow$ 100B Stack (50% Replay) 49.16 43.72 32.60 52.09 50.59 68.93 27.65 46.39
100B DCLM $\rightarrow$ 100B Stack $\rightarrow$ 100B German 21.42 30.27 26.80 51.22 35.60 57.56 27.96 34.66
100B DCLM $\rightarrow$ 100B Stack $\rightarrow$ 100B German (50% Replay) 44.82 40.47 30.6 52.09 48.65 67.35 25.49 44.21
@@
100B DCLM $\rightarrow$ 100B Stack $\rightarrow$ 100B German 19.61 30.01 24.6 49.80 35.82 57.12 28.26 34.21
100B DCLM $\rightarrow$ 100B Stack $\rightarrow$ 100B German (50% Replay) 44.05 41.08 32.8 52.88 48.94 68.22 27.03 45.0
@@
Table 3: Zero-shot results on popular LM benchmarks. Normalized accuracy is reported. The best performing results (Avg.) are bolded for task DCLM $\rightarrow$ Stack. Similarly, for task DCLM $\rightarrow$ Stack $\rightarrow$ German, best performing results are underlined. On average, as expected, we observe that the infinite schedule with $\eta_{const}=2e-4$ with a $50\%$ replay gives the best performance across all the benchmarks. Even without Replay, both the infinite schedules give better performance as compared to cosine. This demonstrates the effectiveness of the infinite schedule in mitigating forgetting.
@@ 7 Conclusion
Our results suggest that infinite cosine schedules offer a flexible and robust framework for continual pre-training (CPT) of foundation models across vision and language domains. They enable seamless training continuation from intermediate checkpoints (e.g., at $\eta_{const}$), support dynamic adaptation strategies such as adjustable replay, and maintain strong performance under distribution shifts, without requiring a predefined training budget. We saw that infinite schedules allow us to seamlessly resume training (continued pre-training at $\eta_{const}$), that alone or along with replay, they best continual learning baselines, and, on large-scale experiments across multiple vision and language datasets, they consistently outperform repeated cosine decay. While we do not claim universal superiority, our large-scale experiments demonstrate that infinite schedules consistently provide competitive retention of prior knowledge and improved stability in non-IID continual learning scenarios, making them a practical alternative to repeated cosine decay in real-world CPT pipelines.
@@
We acknowledge support from NSERC Discovery Grant RGPIN- 2021-04104 [E.B.], FRQNT New Scholar [E.B.], the Canada CIFAR AI Chair Program [I.R.], and the Canada Excellence Research Chairs Program [I.R.]. We would also like to acknowledge funding from the FRQNT Doctoral (B2X) scholarship [B.T.]. This research was made possible thanks to the computing resources on the Frontier supercomputer, provided as a part of the ALCC 2024 program award “Scalable Foundation Models for Transferable Generalist AI”. These resources were provided by the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. In particular, we thank Jens Glaser for his help with the Summit supercomputer. We would also like to thank Isaac and Nafi of ORNL for their assistance in obtaining the MRI dataset used in additional experiments.
@@ A.2 Hyperparameters and Implementation details for CIFAR10 MAE
For our architecture, we employ a ViT-tiny encoder (12 layers, 192 hidden dimensions, 3 attention heads) to match the scale of CIFAR-10, with our implementation based on Zhang (2021)’s work. Our model uses a masking ratio of 0.75, consistent with the original MAE, and incorporates a lightweight decoder (4 layers) with learned position embeddings to reconstruct the masked patches. Regarding the learning rate configuration, we selected a maximum learning rate of 7.5e-5 through hyperparameter tuning over the values [7e-5, 1.5e-4, 3e-4] on the first two tasks, with a minimum learning rate of 7.5e-6. For most experiments, we employ a constant learning rate of 1.875e-5 and a cooldown proportion of 0.4, except for experiments without replay, where we increase the constant learning rate to 5.625e-5. These optimal values were determined through experiments similar to our large-scale setup, testing cooldown proportions [0.3, 0.4, 0.5] and constant learning rates [1.875e-5, 5.625e-5]. While these findings align with our large-scale experiments, the small dataset size necessitated a slightly larger cooldown proportion to maintain a higher learning rate for a longer duration. For linear probing experiments in our small-scale setup, we utilized the AdamW (Loshchilov & Hutter, 2019) optimizer with a weight decay coefficient of 5e35e-3 and momentum parameters $\beta_{1}$ and $\beta_{2}$ set to 0.9 and 0.95, respectively. The linear probing experiments implemented a cosine decay learning rate schedule with a maximum learning rate $\eta_{max}=1e-3$, running for 100 epochs total, including 10 warmup epochs, with a batch size of 128. Complete hyperparameter details for pre-training and linear probing can be found in the corresponding tables Table 6 and Table 6. For the baseline methods MAS (Aljundi et al., 2018) and LwF (Li & Hoiem, 2017), we conducted hyperparameter tuning using grid search over the first two tasks. For MAS, we explored values of $\alpha$ and $\lambda$ in [0.25, 0.5, 0.75]. Similarly for LwF, we searched for optimal $\alpha$ values within the same range.
@@ Appendix C Implementation details and hyperparameters for language pre-training
Warmup percent ($N_{\textit{w}}$) 11
@@
Warmup percent ($N_{\textit{w}}$) 11
Cooldown iters percent ($N_{\textit{c}}$) 0
@@ Appendix D Pretraining with DCLM data
Figure 7 shows the validation loss on the DCLM dataset for cosine and infinite schedule with varying $\eta_{const}$ and cooldown proportion $P$. We observe that the infinite schedule with a higher constant learning rate ($\eta_{const}=2e-4$) and cooldown proportion ($P=0.6$) performs better than the cosine schedule and the other configurations of the infinite schedule. The final checkpoint, in the case of infinite schedule, is obtained via annealing which we perform for $15\%$ of the total iterations after the constant phase, as shown in Figure 1. It can be inferred that the infinite schedule with $\eta_{const}=2e-5$ and $P=0.6$ performs the best, with validation loss rapidly decaying in the annealing phase. We also observe that a shorter cooldown phase ($P=0.3$) results in suboptimal performance with higher validation loss, thus indicating that a longer cooldown phase is beneficial. We note that this corresponds to $28K$ steps. As for the $\eta_{const}$, we observe that both 1e41e-4 and 2e42e-4 perform similarly, with the latter having a slightly lower validation loss, indicating that a higher constant learning rate gives a better exploration possibility during training.
@@ Appendix E Pre-training with combined DCLM and Stack Data
(a) Valid. Loss on DCLM
(b) Valid. Loss on Stack
(c) Valid. Loss on German
Figure 9: Validation Loss ($\downarrow$ is better) for different schedules accompanied with replay. The total fraction of replay is 50%, with 25% of DCLM and other 25% from Stack Data. CPT is on German data ($\mathcal{D}_{2}$), validating on all German ($\mathcal{D}_{2}$) DCLM ($\mathcal{D}_{0}$) and Stack ($\mathcal{D}_{1}$) datasets. We further perform an experiment with equal proportion of data during annealing, referred to as Eq. in the above graphs. It can be observed that equal proportions improve the upstream performance (lower validation loss on DCLM and Stack). The downstream performance is quite similar to other Infinite schedules (both $\eta_{const}\in\{1e-4,2e-4\}$) config. The downstream performance of the infinite schedule on the current task (German) is comparable to cosine.
Appendix F Dynamic Adjustment of Replay Buffer with equal proportion of pre-training data during Annealing Phase
We perform an ablation by annealing on equal proportion of data, i.e. 33.33% each of DCLM ($\mathcal{D}_{0}$), Stack ($\mathcal{D}_{1}$) and German ($\mathcal{D}_{2}$). As shown in Figure 9, it can be inferred that the performance on previous tasks improve as compared to the config where we use 50% of buffer, since now the proportion of previous tasks data has increased (25% for each of DCLM and Stack $\rightarrow$ 33.33%). But this does not deteriorate the downstream performance on German data. Hence in cases, where upstream performance is critical, we can anneal on an equal proportion of data.
Appendix G Effect of Checkpoint selection in Past task performance
Figure 10: Comparison of linear probe loss on Imagenet ($D_{0}$) with checkpoints obtained while training Places 2 ($D_{1}$). All experiments were conducted with replay. We compare different checkpoint selection strategies against the infinite scheduler. Infinite Cosine learning rate schedule (green) achieves consistently lower loss compared to cosine schedules restarted from either checkpoint at $\eta_{const}$ (yellow) or $N_{d}$ (blue).
To address the relative importance of the learning rate scheduling function versus the choice of checkpoint, we conducted an additional ablation study. Figure 10 shows the performance comparison of three different approaches:
  1. 1.
    Cosine learning rate schedule with checkpoint from $\eta_{\text{const}}$ checkpoint (yellow line): This checkpoint is obtained at the point where the cosine schedule reaches the learning rate corresponding to the value of $\eta_{\text{const}}$ used in our infinite cosine schedule.
  2. 2.
    Cosine learning rate schedule with checkpoint from $N_{d}$ checkpoint (blue line): This checkpoint is obtained at the time corresponding to when the annealing of the infinite schedule begins.
  3. 3.
    Infinite Cosine learning rate schedule with $\eta_{\text{const}}=3.75\text{e-}5$ and $P=0.3$ (green line): Our best performing scheduler from main paper.
As shown in Figure 10, the Infinite Learning Rate schedule consistently achieves lower linear probe loss throughout training compared to both cosine schedule variants. The results demonstrate that this approach provides more effective representation learning compared to cosine schedule with early checkpointing.
Strategic checkpoint selection plays a critical role in preserving previously acquired knowledge. However, this flexibility is not inherently supported by the standard cosine schedule, which is typically designed to run uninterrupted until the end of training. In our comparison experiments, we manually extracted checkpoints at arbitrary points with high learning rates which is an approach that deviates from standard usage of cosine schedule. In contrast, the Infinite Learning Rate schedule naturally enables such flexibility through its constant learning rate phase, explicitly designed for this requirement. This built-in mechanism aids in knowledge retention across tasks, contributing to the consistently lower linear probe loss observed compared to other approaches.
Appendix H Additional Forgetting metrics
We present a more thorough analysis of forgetting dynamics in our continual learning framework in Table 11. While our main paper reports standard metrics which measured at the end of the training including Average Accuracy (AA), Forward Transfer (FWT), and Backward Transfer (BWT), this section extends our evaluation with additional metrics proposed by D´ıaz-Rodr´ıguez et al. (2018) to provide a more complete picture of the retention capabilities of our approach after every task.
D´ıaz-Rodr´ıguez et al. (2018) provides metrics to analyze the performance of the model at every timestep to incorporate the dynamic nature of CL. Hence we use an overall prefix for these metrics which are taken as an average after every time step. Overall Accuracy ($A=\frac{\sum_{i\geq j}^{T}R_{i,j}}{\frac{T(T+1)}{2}}$) gives the average of accuracy on all tasks after every timestep. Overall Backward transfer ($BWT=\frac{\sum_{i=2}^{T}\sum_{j=1}^{i-1}\left(R_{i,j}-R_{j,i}\right)}{\frac{T(T-1)}{2}}$) measures the backward transfer after every timestep, while Overall Forward transfer ($FWT=\frac{\sum_{i<j}^{T}R_{i,j}}{\frac{T(T-1)}{2}}$) measures the forward transfer after every time step. Overall remembering (REM=100|min(BWT,0)|REM=100-|min(BWT,0)|) quantifies the amount of knowledge remembered by the model across all time steps. Finally, we calculate the CL score to get a weighted average of all the metrics $CL_{score}=\sum_{i=1}^{\#c}w_{i}c_{i}$, where #c\#c denotes the total number of metrics used. We use an equal weighting and report the equal weighted average. We additionally show the performance of a model which is pre-trained on the combination of all datasets, then linear probed on each dataset and obtained accuracy. We report the overall accuracy of it across all tasks as a reference in the table.
Strategy Overall Acc Overall REM Overall BWT Overall FWT CL_score
Cosine + ER (5%) 44.76 97.11 -2.89 12.85 37.96
Infinite cosine + ER (5%) 50.56 99.54 -0.46 12.78 40.61
Cosine 43.61 84.61 -15.39 12.80 31.41
Infinite cosine 44.36 86.98 -13.02 13.08 32.85
Full baseline 53.40 - - - -
Table 11: Comparison of forgetting metrics (as defined in D´ıaz-Rodr´ıguez et al. (2018)) across various methods. Infinite cosine with replay shows competitive results with lower forgetting, approaching the performance of the Full baseline.
Appendix I Practical Guide for Hyperparameter Selection in Infinite Cosine Scheduling
Selecting learning rate hyperparameters for continual pre-training can be computationally expensive due to the scale of training. This challenge applies to both repeated cosine and infinite cosine schedules and is not unique to the latter. However, when control over the initial pre-training phase is available, recent advances in hyperparameter-transfer techniques Yang et al. (2021) can help reduce tuning costs substantially.
While hyperparameter stability is not the primary focus of this work, we provide practical guidance based on consistent trends observed in our experiments. The following empirically grounded rule-of-thumb can serve as a starting point for configuring the infinite cosine schedule effectively:
  • Step 1: Selecting $\eta_{max}$ - Choose a maximum learning rate ($\eta_{max}$) that yields stable validation loss on the initial domain. This follows standard large-scale pre-training practice. If a tuned cosine schedule is already available from prior work, its $\eta_{max}$ serves as a strong candidate.
  • Step 2: Setting $\eta_{const}$ - Define the constant learning rate for Inf-Cos as the midpoint between $\eta_{max}$ and $\eta_{min}$. For example, with $\eta_{max}=3e-4$ and $\eta_{min}=3e-5$, a suitable baseline is $\eta_{\text{const}}\approx 1.65\times 10^{-4}$.
We further find the following trends to be consistent along with takeaways from Hägele et al. (2024)
  • Without replay: When no replay buffer is available, a relatively lower $\eta_{const}$ is preferred to better preserve previously acquired knowledge as shown in Figure 3 and Figure 5.
  • With replay: When replay is available, a relatively higher $\eta_{const}$ can be utilized to improve adaptability without severely impacting retention — as shown in Figure 4.
These insights allow practitioners to configure Infinite Schedules without exhaustive search, and makes it a more flexible and robust scheduler as compared to repeated cosine.
Appendix J Additional Experiments: Continual Learning with Medical Imaging
To evaluate the robustness of our MAE continual pretraining approach under extreme domain shifts, we conduct additional experiments incorporating a fourth task that introduces a substantial distribution change. Specifically, we transition from natural image datasets to specialized medical imaging data, presenting a challenging scenario for knowledge retention and adaptation.
In this extended evaluation, we implement a continual learning scenario with an unlimited buffer capacity while constraining the replay mechanism through batch composition. During training on the fourth task, we enforce a balanced sampling strategy where 50% of each batch consists of samples from the current task (MRI data) and the remaining 50% comprises samples uniformly drawn from the combined buffer containing all previous datasets (ImageNet, Places2, and FireRisk).
The medical imaging dataset consists of multi-planar brain MRI scans. The dataset includes four distinct classes:“Empty” (background tissue),“Whole Tumor”, “Tumor Core”, and “Enhancing Tumor”. For evaluation, we compute the average prediction accuracy across all relevant binary classification from the multi-class labels.
Table 12 presents the linear probe accuracies across all four tasks, comparing our proposed infinite learning rate schedule against the standard cosine annealing baseline. The infinite learning rate schedule consistently outperforms cosine annealing in mitigating catastrophic forgetting, maintaining higher accuracies on ImageNet (57.20% vs 54.11%), Places2 (47.16% vs 46.46%), and FireRisk (60.56% vs 60.32%) after completing all tasks. These results validate the effectiveness of our approach in continual learning scenarios involving substantial domain shifts.
Task completed Acc on Imagenet Acc on Places Acc on FireRisk Acc on MRI
Cosine Infinite Cosine Infinite Cosine Infinite Cosine Infinite
After Imagenet 60.72 59.84 46.11 45.70 59.80 59.89 85.49 85.46
After Places 58.89 60.35 48.71 48.37 60.24 60.73 85.50 85.46
After Firerisk 54.35 57.13 45.98 47.23 61.06 61.29 85.48 85.47
After MRI 54.11 57.20 46.46 47.16 60.32 60.56 85.93 85.92
Table 12: Linear probe accuracy comparison across four continual learning tasks using 50% buffer replay. The infinite learning rate schedule consistently outperforms cosine annealing in mitigating catastrophic forgetting across diverse domains. Gray cells indicate forward transfer evaluation (tasks not yet encountered).